three bikers,allan,anton, and alex starts at the same spot.allan can complete a whole round in 50 mins,anton in 48 mins,and alex in 36 mins.if they all start at 3 in the afternoon at what time will they be on the same spot again?
Answer by ankor@dixie-net.com(22740)

three bikers,allan,anton, and alex starts at the same spot.
allan can complete a whole round in 50 mins,
anton in 48 mins,and
alex in 36 mins.
If they all start at 3 in the afternoon at what time will they be on the same spot again?
:
Find the least common multiple of 50, 48, 36
Prime factors
50: 2 * 5 * 5
48: 2 * 2 * 2 * 2 * 3
36: 2 * 2 * 3 * 3
:
LCM = 2 * 2 * 2 * 2 * 3 * 3 * 5 * 5 = 3600 min, 3600/60 = 60 hrs which is 2.5 days
The will all be at the same spot at 3 AM (2.5 days later)